# TrialBench: Multi-Modal AI-Ready Clinical Trial Datasets

### v2

@@ 3.1 Data Overview
  • Phase. Phase I tests the toxicity and side effects of the drug; phase II determines the efficacy of the drug (i.e. if the drug works); phase III focuses on the effectiveness of the drug (i.e., whether the drug is better than the current standard practice). When the trial passes phase III, it can be submitted to the FDA for approval. In many cases, even after approval, we still need to further monitor the drugs’ effectiveness and safety. Sometimes a phase IV trial will be conducted to assess the drug’s effectiveness and safety. Table 3 demonstrate the differences between phases I, II, III, and IV.
  • Eligibility criteria describe the patient recruitment requirements in unstructured natural language. Eligibility criteria comprise multiple inclusion and exclusion criteria, which specify what is desired and undesired when recruiting patients. Each individual criterion is usually a natural language sentence. For example, in the clinical trial entitled “Efficacy and Safety Study of MP-513 in Combination With Thiazolidinedione in Patients With Type 2 Diabetes ”fffhttps://clinicaltrials.gov/ct2/show/NCT01026194, which is a phase III trial, the inclusion criteria contain:
@@
  • Sponsors of the clinical trial can be pharmaceutical companies or research institutes. For example, the trial entitled “PF-06863135 As Single Agent And In Combination With Immunomodulatory Agents In Relapse/Refractory Multiple Myeloma”ggghttps://clinicaltrials.gov/ct2/show/NCT03269136 is supported by Pfizer; the trial entitled “Five, Plus Nuts and Beans for Kidneys” hhhhttps://clinicaltrials.gov/ct2/show/NCT03299816 is supported by Johns Hopkins University. Some trials may contain multiple sponsors. Table 4 lists the top 20 sponsors that conduct the most interventional clinical trials.
@@
Table 5: A real example of a clinical trial record.
Feature Descriptions
NCTID NCT00610792
disease Ovarian Cancer
phase II
study type interventional
drug bortezomib and pegylated liposomal doxorubicin
start date July 2006
completed date September 2009
sponsor Millennium Pharmaceuticals, Inc.
outcome withdrawn
3.2 Summarization of Multi-Modal Features
@@
In clinical trials, there are many text features that contain rich information for artificial intelligence modeling. For example, eligibility criteria describe the patient recruitment requirements in unstructured natural language; each clinical trial contains a summary, which consists of 2-5 natural language sentences that describe the tested treatment, the target disease to treat, and the main objective of the clinical trial. To process such datasets, we treat the text data as sequences of tokens (e.g., words). How to extract useful information from unstructured text has been extensively studied with several well-known deep neural network architectures, such as recurrent neural network (RNN) [50], convolutional neural network (CNN), and transformer architecture [51].
@@
Diseases lie at the heart of clinical trial and drug discovery. There are several standardized disease coding systems that healthcare providers use for the electronic exchange of clinical health information, including the International Classification of Diseases, Tenth Revision, Clinical Modification (ICD-10-CM), The International Classification of Diseases, Ninth Revision, Clinical Modification (ICD-9-CM), and Systematized Nomenclature of Medicine – Clinical Terms (SNOMED CT) [53]. These coding systems contain disease concepts organized into hierarchies. We take the ICD-10-CM code as an example. ICD-10-CM is a seven-character, alphanumeric code. Each code begins with a letter, and two numbers follow that letter. The first three characters of ICD-10-CM are the “category”. The category describes the general type of injury or disease. A decimal point and the subcategory follow the category. For example, the code “G44” represents “Other headache syndromes”; the code “G44.31” represents “Acute post-traumatic headache”; the code “G44.311” represents “Acute post-traumatic headache, intractable”. G44.311 has two ancestors: G44 and G44.31, where an ancestor represents a higher-level category of the current code. The description of all the ICD-10-CM codes is available at https://www.icd10data.com/ICD10CM/Codes. We also illustrate the hierarchy in Figure 3.
@@
4 Technical Validation
To show the processed datasets are AI-ready and of reasonable quality, we evaluate the performance of these datasets on mainstream artificial intelligence algorithms. We leverage a multi-modal deep neural network to represent the multi-modal features and concatenate all these representations to make the prediction. In this section, we first discuss the multi-modal deep learning method, then describe the experimental setup, and present the experimental results finally.
4.1 Multi-modal Deep Neural Networks
For all the classification and regression tasks, we apply various deep neural networks to represent multi-modal features. Each representation is an embedding vector with continuous values. Then, we concatenate these representations, feed them into multiple layer perceptron (MLP), and make the prediction. For the eligibility criteria design task, we use OpenAI ChatGPT APIiiihttps://openai.com/index/openai-api/ with the prompt to produce eligibility criteria.
@@ 4.1 Multi-modal Deep Neural Networks
Recently, numerous tabular data processing models [48, 54, 49] have been proposed for numerical and categorical feature processing. Among them, DANets [11] stand out due to its key component’s modularity and ability to achieve competitive performance without hyper-parameter tuning. The key component, the basic block module, supports flexible stacking, making DANets suitable as a submodule for processing numerical and categorical features. After preprocessing (e.g., normalization), three lightweight basic blocks are sequentially stacked to hierarchically select, extract, and merge features from input categorical and numerical features, ultimately yielding a 50-dimensional embedding.
@@
After obtaining the representations of multi-modal data, we concatenate these representations, feed the concatenated vector into multiple-layer perceptron (MLP), and make the prediction. For binary classification tasks (e.g., trial approval prediction), we use sigmoid function as the activation function in the output layer to yield predicted probability; for multi-category classification tasks (e.g., trial failure reason identification), we use softmax as the activation function in the output layer to produce probability distribution over all the categories; for regression tasks (e.g., trial duration prediction), we do not use activation function in the output layer to produce continuous-valued prediction. We use cross-entropy criterion as the loss function for classification tasks and mean-square error (MSE) as the loss function for regression tasks.
@@ 4.2 Experimental Setup
All the code is implemented in Python 3.8. All the deep learning models are implemented in PyTorch, and we use GPT 4.0 for data annotation and the generation tasks. The embedding size of all the representations is set to 100. We use Adam [60] as the numerical optimizer to minimize the loss function with an initial learning rate at $1e{-3}$ and zero weight decay. The batch size is set to 64. The maximal training epochs is set to 20.
@@
For classification tasks, we assess the model performance using accuracy, PR-AUC (the area under the Precision-Recall curve), F1 score (the harmonic mean of precision and recall), and ROC-AUC (the Area Under the Receiver Operating Characteristic Curve). For regression tasks, we use RMSE (Root Mean Squared Error), MAE (Mean Absolute Error), Concordance Index, and Pearson Correlation as metrics. For generation tasks (eligibility criteria design), we design some semantic metrics to measure the alignment between real and designed criteria, including text embeddings’ cosine similarity, informativeness, and redundancy, detailed in the Appendix.
4.3 Validation Results
@@ 5 Usage Note
AI4Trial is intended for healthcare, biomedical, and artificial intelligence researchers and data scientists who want to apply AI algorithms and innovate novel methods to tackle problems formulated in TrialBench datasets and tasks.
@@
The curated dataset are available at https://huggingface.co/datasets/ML2Healthcare/ClinicalTrialDataset and the validation frameworks are available at https://github.com/ML2Health/ML2ClinicalTrials/tree/main/AI4Trial.
@@
Appendix A Evaluation Metrics
Classification Metrics
Binary classification is the most widely used classification. For example, in clinical trial approval prediction, there are usually binary labels, 1 denotes the trial passes while 0 indicates fail. The prediction can be the probability that the trial pass. In binary classification, there are four kinds of test data points based on their ground truth and the model’s prediction,
  1. 1.
    positive sample and is correctly predicted as positive, also known as True Positive (TP);
  2. 2.
    negative samples and is wrongly predicted as positive samples, also known as False Positive (FP);
  3. 3.
    negative samples and is correctly predicted as negative samples, also known as True Negative (TN);
  4. 4.
    positive samples and is wrongly predicted as negative samples, also known as False Negative (FN).
We use different evaluation metrics for multi-category classification and binary classification.
  • Precision. The precision is the performance of a classifier on the samples that are predicted as positive. It is formally defined as $\text{precision}=\frac{TP}{TP+FP}$.
  • Recall. The recall score measures the performance of the classifier to find all the positive samples. It is formally defined as $\text{recall}=\frac{TP}{TP+FN}$.
  • Accuracy. Accuracy is the fraction of correctly predicted/classified samples. It is formally defined as $\text{accuracy}=\frac{TP+TN}{TP+TN+FP+FN}$. It is worth mentioning that when the data samples are highly imbalanced, e.g., positive samples are far more than negative ones, the accuracy is not a proper metric as it is trivial to achieve high accuracy by always predicting the majority class.
  • PR-AUC (Precision-Recall Area Under Curve). The area under the Precision-Recall curve summarizes the trade-off between the true positive rate and the positive predictive value for a predictive model using different probability thresholds.
[…]
  • ROC-AUC Area Under the Receiver Operating Characteristic Curve summarizes the trade-off between the true positive rate and the false positive rate for a predictive model using different probability thresholds. ROC-AUC is also known as Area Under the Receiver Operating Characteristic curve (AUROC) in some literature.
For all these metrics, the numerical values range from 0 to 1, a higher value represents better performance. We typically report multiple metrics to measure the performance comprehensively.
Regression Metrics
In the regression task, both ground truth and prediction are continuous values.
  • Mean Squared Error (MSE) measures the average of the squares of the difference between the forecasted value and the actual value. It is defined as $\text{MSE}=\frac{1}{N}\sum^{N}_{i=1}(y_{i}-{\widehat{y}}_{i})^{2}$, where $N$ is the size of the test set; $y_{i}$ and $\widehat{y}_{i}$ denote the ground truth and predicted score of the $i$-th data sample in the test set, respectively. MSE value ranges from 0 to positive infinity. A lower MSE value indicates better performance.
  • Mean Absolute Error (MAE) measures the absolute value of the difference between the predicted value and the actual value. It is defined as $\text{MAE}=\frac{1}{N}\sum^{N}_{i=1}|y_{i}-{\widehat{y}}_{i}|$, where $N$ is the size of the test set; $y_{i}$ and $\widehat{y}_{i}$ denote the ground truth and predicted score of the $i$-th data sample in the test set, respectively. MAE value ranges from 0 to positive infinity. It emphasizes the ranking order of the prediction instead of the absolute value. A lower MAE value indicates better performance.
  • Concordance Index (also known as c-index) is defined as the proportion of concordant pairs divided by the total number of possible evaluation pairs. A higher value represents better prediction performance.
  • Pearson Correlation (PC) is defined as the covariance of the prediction and the ground truth divided by the product of their standard deviations. For two random variables $x$ and $y$, Pearson Correlation is formally defined as
    $\text{PC}=\frac{\mathbb{E}[(x-\mu_{x})({y}-\mu_{y})]}{\sigma_{x}\sigma_{y}},$ (5)
    In the regression task, suppose there are $N$ data points in the test set, $y_{i}$ is the ground truth of the $i$-th data sample, $\widehat{y}_{i}$ is the prediction for $i$-th data, Pearson Correlation becomes
    $\text{PC}=\frac{\sum_{i=1}^{N}\big{(}(y_{i}-\mu_{y})(\widehat{y}_{i}-\mu_{\widehat{y}})\big{)}}{\sigma_{y}\sigma_{\widehat{y}}},$ (6)
    where $\mu_{y}=\frac{1}{N}\sum_{j=1}^{N}y_{j}$ and $\mu_{\widehat{y}}=\frac{1}{N}\sum_{j=1}^{N}\widehat{y}_{j}$ are mean of ground truth and prediction, respectively. $\sigma_{y}=\sum_{i=1}^{N}(y_{i}-\frac{1}{N}\sum_{j=1}^{N}y_{j})^{2}$ and $\sigma_{\widehat{y}}=\sum_{i=1}^{N}(\widehat{y}_{i}-\frac{1}{N}\sum_{j=1}^{N}\widehat{y}_{j})^{2}$ are the standard deviations of ground truth and prediction, respectively. The value ranges from -1 to 1. A higher Pearson correlation value indicates better performance.
  • R-squared ($R^{2}$) score is defined as the proportion of the variation in the dependent variable that is predictable from the independent variable(s). It is also known as the coefficient of determination in statistics. Suppose we have $N$ continuous ground truth $y_{1},\cdots,y_{N}$ and $N$ corresponding predictions $\widehat{y_{1}},\cdots,\widehat{y_{N}}$. The difference $y_{i}-\widehat{y_{i}}$ is called residual. Then, we defined the residual sum of squares as
    $\text{SS}_{\text{res}}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\widehat{y_{i}})^{2},$ (7)
    and define the total sum of squares as
    $\text{SS}_{\text{total}}=\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\bar{y})^{2},$ (8)
    where $\bar{y}=\frac{1}{N}y_{i}$ is the mean of the ground truth. Then $R^{2}$ score is defined as
    $R^{2}=1-\frac{\text{SS}_{\text{res}}}{\text{SS}_{\text{total}}}.$ (9)
    Higher $R^{2}$ scores indicate better performance. In a perfect prediction model, the prediction exactly matches the ground truth, then $\text{SS}_{\text{res}}=0$ and $R^{2}=1$. A weak predictor that always predicts $\bar{y}$ would cause $R^{2}=0$. Some predictors would even lead to a negative $R^{2}$ score.
  • Spearman’s rank correlation coefficient is the Pearson correlation coefficient between the rank variables. Given two groups of variables $X$ and $Y$, it is formally defined as
    Spearman $=\text{PC}(R(X),R(Y))=\frac{\mathbb{E}\big{[}(R(X)-\mu_{R(X)})(R(Y)-\mu_{R(Y)})\big{]}}{\sigma_{R(X)}\sigma_{R(Y)}},$ (10)
    where PC is Pearson Correlation defined above; $R(X)$ is the rank variable of $X$, whose value is from the set $\{1,2,\cdots,|R(X)|\}$. In regression tasks, $X$ and $Y$ represent the ground truth and prediction, respectively.
In some cases, we are interested in the difference between ground truth and prediction, and would typically choose mean squared error or mean absolute error as the evaluation metrics. Sometimes, we are interested in the correlation between ground truth and prediction. In these cases, we prefer to use Pearson correlation and R-squared score ($R^{2}$ score). Also, in some applications, the rank-ordering of predictions is more important than the difference between prediction and ground truth. In this case, we prefer to use Spearman’s rank correlation coefficient or concordance index.
Generation Metrics
In the generation tasks (e.g., eligibility criteria design), we design a couple of metrics to evaluate the quality of the generated text content.
  • Cosine Similarity between text embeddings. To measure the semantic similarity, we use OpenAI API (GPT 4.0) jjjhttps://openai.com/index/openai-api/ to obtain the text embeddings of generated and groundtruth content and then calculate their cosine similarity.
  • Informativeness. In text generation, informativeness is one of the key metrics for evaluating the quality of generated text. Informativeness measures the richness of useful information provided by the generated text. High informativeness usually means the text contains more relevant details, facts, and insights, rather than simply repeating known information or containing empty content. We use TF-IDF (Term Frequency-Inverse Document Frequency, a well-known metric in natural language processing) to measure the informativeness of the generated text relative to a reference text: (1) TF-IDF vectorization: Combine the generated text and reference text, and vectorize them using the TF-IDF method. TF-IDF is a statistical method used to evaluate the importance of a word in a document. (2) cosine similarity: Calculate the cosine similarity between the generated text’s TF-IDF vector and the reference text’s average TF-IDF vector. The higher the cosine similarity, the more similar the generated text is to the reference text. (3) informativeness score: The informativeness score is calculated as 1 minus cosine similarity. A higher score indicates that the generated text contains more different information relative to the reference text, thus having greater informativeness.
  • Redundancy. Redundancy refers to the presence of unnecessary repetitive information or extraneous content in the text, which can affect the quality and readability of the text. Methods for evaluating redundancy typically involve detecting repeated words, phrases, or sentences. We need the following four steps for measuring the redundancy of generated text: (1) N-gram list: Split the generated text into a list of words and generate an N-gram list. (2) count N-gram occurrences: Count the number of times each N-gram appears in the text. (3) calculate unique N-gram quantity: Count the number of N-grams that appear only once. (4) calculate redundancy: Redundancy is calculated as 1 minus the ratio of the number of unique N-grams to the total number of N-grams.
Figure 5: Hierarchical features presented in XML for each registered clinical trial.
Figure 6: Input Variables for trial duration prediction.
Figure 7: Features for patient dropout prediction, serious adverse event prediction, and mortality event prediction.
Figure 8: Feature for trial approval prediction.
Figure 9: Feature for trial failure reason identification.
Figure 10: Feature for eligibility criteria design.

### v3

@@ Methods
  • Phase. Phase I tests the toxicity and side effects of the drug; phase II determines the efficacy of the drug (i.e. if the drug works); phase III focuses on the efficacy of the drug (i.e., whether the drug is better than the current standard practice). When the trial passes phase III, it can be submitted to the FDA for approval. In many cases, even after approval, we still need to further monitor the drugs’ efficacy and safety. Sometimes a phase IV trial will be conducted to assess the drug’s efficacy and safety. Table 4 demonstrate the differences between phases I, II, III, and IV.
  • Eligibility criteria describe the patient recruitment requirements in unstructured natural language. Eligibility criteria comprise multiple inclusion and exclusion criteria, which specify what is desired and undesired when recruiting patients. Each individual criterion is usually a natural language sentence. For example, in the clinical trial entitled “Efficacy and Safety Study of MP-513 in Combination With Thiazolidinedione in Patients With Type 2 Diabetes ”(https://clinicaltrials.gov/ct2/show/NCT01026194), which is a phase III trial, the inclusion criteria contain:
@@
  • Sponsors of the clinical trial can be pharmaceutical companies or research institutes. For example, the trial entitled “PF-06863135 As Single Agent And In Combination With Immunomodulatory Agents In Relapse/Refractory Multiple Myeloma”(https://clinicaltrials.gov/ct2/show/NCT03269136) is supported by Pfizer; the trial entitled “Five, Plus Nuts and Beans for Kidneys” (https://clinicaltrials.gov/ct2/show/NCT03299816) is supported by Johns Hopkins University. Some trials may contain multiple sponsors. Table 5 lists the top 20 sponsors that conduct the most interventional clinical trials.
@@
Summarization of Multi-Modal Features
@@
In clinical trials, there are many text features that contain rich information for AI modeling. For example, eligibility criteria describe the patient recruitment requirements in unstructured natural language; each clinical trial contains a summary, which consists of 2-5 natural language sentences that describe the tested treatment, the target disease to treat, and the main objective of the clinical trial. To process such datasets, we treat the text data as sequences of tokens (e.g., words). How to extract useful information from unstructured text has been extensively studied with several well-known deep neural network architectures, such as recurrent neural network (RNN) [44], convolutional neural network (CNN), and transformer architecture [45].
@@
There are several standardized disease coding systems that healthcare providers use for the electronic exchange of clinical health information, including the International Classification of Diseases, Tenth Revision, Clinical Modification (ICD-10-CM), The International Classification of Diseases, Ninth Revision, Clinical Modification (ICD-9-CM), and Systematized Nomenclature of Medicine – Clinical Terms (SNOMED CT) [47]. These coding systems contain disease concepts organized into hierarchies. We take the ICD-10-CM code as an example. ICD-10-CM is a seven-character, alphanumeric code. Each code begins with a letter, and two numbers follow that letter. The first three characters of ICD-10-CM are the “category”. The category describes the general type of injury or disease. A decimal point and the subcategory follow the category. For example, the code “G44” represents “Other headache syndromes”; the code “G44.31” represents “Acute post-traumatic headache”; the code “G44.311” represents “Acute post-traumatic headache, intractable”. G44.311 has two ancestors: G44 and G44.31, where an ancestor represents a higher-level category of the current code. The description of all the ICD-10-CM codes is available at https://www.icd10data.com/ICD10CM/Codes. We also illustrate the hierarchy in Figure 3.
@@
Technical Validation
To show the processed datasets are AI-ready and of reasonable quality, we evaluate the performance of these datasets on mainstream AI algorithms. We leverage a multi-modal deep neural network to represent the multi-modal features and concatenate all these representations to make the prediction. In this section, we first discuss the multi-modal deep learning method, then describe the experimental setup, and present the experimental results finally.
Multi-modal Deep Neural Networks
For all classification and regression tasks, we apply various deep neural networks to represent multimodal features. Each representation is an embedding vector with continuous values. Then, we concatenate these representations, feed them into multiple layer perceptron (MLP), and make the prediction. For the eligibility criteria design task, we use OpenAI ChatGPT API(https://openai.com/index/openai-api/) with the prompt to produce eligibility criteria.
@@
Recently, numerous tabular data processing models [42, 48, 43] have been proposed for numerical and categorical feature processing. Among them, DANets [16] stand out due to its key component’s modularity and ability to achieve competitive performance without hyperparameter tuning. The key component, the basic block module, supports flexible stacking, making DANets suitable as a submodule for processing numerical and categorical features. After preprocessing (e.g., normalization), three lightweight basic blocks are sequentially stacked to hierarchically select, extract, and merge features from input categorical and numerical features, ultimately yielding a 50-dimensional embedding.
@@
After obtaining the representations of multi-modal data, we concatenate these representations, feed the concatenated vector into a multiple-layer perceptron (MLP), and make the prediction. For binary classification tasks (e.g., trial approval prediction), we use the sigmoid function as the activation function in the output layer to yield predicted probability; for multi-category classification tasks (e.g., trial failure reason identification), we use softmax as the activation function in the output layer to produce probability distribution over all the categories; for regression tasks (e.g., trial duration prediction), we do not use activation function in the output layer to produce continuous-valued prediction. We use cross-entropy criterion as the loss function for classification tasks and mean-square error (MSE) as the loss function for regression tasks.
@@
All the code is implemented in Python 3.8. All deep learning models are implemented in PyTorch, and we use GPT 4.0 for data annotation and generation tasks. The embedding size of all the representations is set to 100. We use Adam [54] as the numerical optimizer to minimize the loss function with an initial learning rate at $1e{-3}$ and zero weight decay. The batch size is set to 64. The maximal training epochs is set to 20.
@@
For classification tasks, we assess the model performance using accuracy, PR-AUC (the area under the Precision-Recall curve), F1 score (the harmonic mean of precision and recall), and ROC-AUC (the Area Under the Receiver Operating Characteristic Curve). For regression tasks, we use RMSE (Root Mean Squared Error), MAE (Mean Absolute Error), Concordance Index, and Pearson Correlation as metrics. For generation tasks (eligibility criteria design), we design some semantic metrics to measure the alignment between real and designed criteria, including text embeddings’ cosine similarity, informativeness, and redundancy, detailed in the Supplementary Information.
Validation Results
@@
TrialBench is intended for healthcare, biomedical, and AI researchers and data scientists who want to apply AI algorithms and innovate novel methods to tackle problems formulated in TrialBench datasets and tasks.
@@
The curated datasets are available at https://zenodo.org/records/14975339; we also make our own website https://huyjj.github.io/Trialbench/, which contains all the datasets, explanations and diagrams for ease of understanding.
Author Information
Contributions
The project was designed by J. Chen, C. Xiao, J. Sun, L. Glass, M. Zitnik, and T. Fu. J. Chen, Y. Hu, Y. Lu, and Y. Wang curated the datasets. Y. Hu, Y. Lu, Y. Wang, and T. Fu developed and validated the model. J. Chen, X. Cao, K. Huang, and T. Fu drafted the paper, while J. Chen, Y. Lu, M. Lin, H. Xu, J. Wu, K. Huang, and T. Fu reviewed and proofread the manuscript.
@@
Corresponding author
Correspondence to Prof. Tianfan Fu at futianfan@gmail.com.